# Clarify Confused Nodes via Separated Learning

### v3

@@ Abstract
Graph neural networks (GNNs) have achieved remarkable advances in graph-oriented tasks. However, real-world graphs invariably contain a certain proportion of heterophilous nodes, challenging the homophily assumption of classical GNNs and hindering their performance. Most existing studies continue to design generic models with shared weights between heterophilous and homophilous nodes. Despite the incorporation of high-order messages or multi-channel architectures, these efforts often fall short. A minority of studies attempt to train different node groups separately but suffer from inappropriate separation metrics and low efficiency. In this paper, we first propose a new metric, termed Neighborhood Confusion (NC), to facilitate a more reliable separation of nodes. We observe that node groups with different levels of NC values exhibit certain differences in intra-group accuracy and visualized embeddings. These pave the way for Neighborhood Confusion-guided Graph Convolutional Network (NCGCN), in which nodes are grouped by their NC values and accept intra-group weight sharing and message passing. Extensive experiments on both homophilous and heterophilous benchmarks demonstrate that our framework can effectively separate nodes and yield significant performance improvement compared to the latest methods. The source code will be released soon.
@@
Graph-structured data can effectively model real-world interactive systems and has received considerable attention recently. Graph neural networks (GNNs) have been proven to be powerful in many graph-related applications, such as anomaly detection [1, 2, 3], fraud detection [4, 5] and recommendation systems [6, 7, 8]. Classical GNNs [9, 10, 11] are adept at learning node representations in graphs where the vast majority of nodes are homophilous, i.e., nodes are primarily connected to nodes with same labels. However, real-world graphs often contain heterophilous nodes [12] and are known as heterophilous graphs when the proportion of these nodes is high, which challenges the homophily assumption of classical GNNs and resulting in diminished performance [13, 14]. In heterophilous graphs, the patterns of connections between nodes are usually complex and varied, rather than being limited to connections with similarly labeled nodes. This is referred to as the heterophily problem, and recent studies [15, 16] have highlighted that the key to addressing this problem is to focus on the diversity of nodes in the graph. Inspired by this, it’s natural to designing flexible model architectures for characterizing diverse nodes, whether homophilous or heterophilous.
@@
Most recent studies endeavor to design general models that still enable weight sharing between homophilous and heterophilous nodes. However, even with the incorporation of higher-order neighbors [17, 18, 19] or multi-channel architectures [20, 21], these models fail to perform consistently well across all evaluation datasets. This may be due to two reasons: 1) these methods do not fundamentally alter the message-passing mechanism that essentially conforms to homophily assumption, thereby introducing noise [22]; 2) nodes within the same graph may exhibit significant behavioral differences due to distribution bias [23, 24]. For example, categorizing an interdisciplinary paper relies on more specific criteria compared to a paper confined to a single discipline [25].Therefore, sharing weights universally across all nodes in a real-world graph is impractical, prompting the necessity for a metric to separate nodes with different characteristics. While numerous studies have explored the separation of neighbors [26, 27] or spectral signals [20, 21], the separation of target nodes or transformation weights, which we refer to as separated learning, remains under-explored. Only a minority of studies [12, 26] have focused on this separating scheme, but they suffer from incorrect separation metrics and low efficiency.
@@
Beginning with the classical Node Homophily (NH) metric [28] derived directly from the definition of homophily or heterophily, we analyze the experiment results on Graph Convolutional Network (GCN) as depicted in Fig. 2. Unexpectedly, the extremely heterophilous (lowest NH) group shows anomalously high accuracy, suggesting that NH fails to separate nodes appropriately. This phenomenon is also observed by recent studies [29, 30, 20], yet they neither ease this anomaly directly with their models nor explore it on a wider range of real-world graphs. Moreover, analytical papers [31, 32, 33] reach the consensus that message passing with too many layers leads to over-smoothing or indistinguishability on node representations, highlighting the importance of distinguishability over consistency with central nodes for model performance. Therefore, we propose a new metric termed Neighborhood Confusion (NC) and detail its superiority over NH. Our proposed framework is guided by this metric for node separation, further validating the effectiveness of NC. Our contributions can be summarized as follows:
@@ II-A Heterophily Problem
Spatial methods allow for fine-grained adaptations of message passing and transformation rules [36]. Geom-GCN [28] and WRGAT [37] rewire the graphs to aggregate neighbors with structural similarities in latent space. SNGNN [38] and GGCN [27] construct a similarity matrix, drawing upon the embeddings from the preceding layer, to separate the dissimilar neighbors. These methods require computing new topologies or similarities and thus are time and memory consuming.
@@ IV Metrics and Observations
Heterophily is harmful to classical GNNs, as different classes of nodes will be inappropriately mixed during message aggregation, leading to the indistinguishability of node representations [30]. However, GCN still performs well on bipartite graphs, which are completely heterophilous under the definition of NH, bringing paradox [29]. As shown in Fig. 2, means that NH metric considers the nodes in the bipartite graph to be extremely heterophilous. To alleviate this paradox, some new metrics have been proposed, such as aggregated similarity score [20], which is calculated by the similarity of intermediate representations, but with high computational complexity.
@@ IV-A Neighborhood Confusion
where is the set of label classes, normalizes NC value to , and higher value indicates a greater degree of label confusion within target neighborhood. The receptive field of NC can be 1- or 2-hop by setting or , which is adaptively configured across different datasets.
@@
NC metric can outline the label distribution in the -hop ego-net of node , and can be regarded as a loose, non-probabilistic form of Shannon entropy, but with reduced computational complexity through the maximization operation.
@@
The non-probabilistic Shannon entropy of the label distribution sampled around (and include) target node is:
Let , then NC is a loose form of because:
The range of is determined to be based on the Maximum Entropy Principle. Consequently, both and can be normalized within the interval of by dividing them with . However, it should be noted that only the maximum value of is used for calculating NC, facilitating a reduction in computational complexity.
@@ IV-B Observations
Rethinking the scenario of bipartite graph, as depicted in Fig. 2, suggests that our metric perceives the label distribution of node neighborhoods to have a low level of confusion, which is quite different from the result of NH. Then, we will further explain why NC surpasses NH in separating nodes from two observations. As an illustrative analysis, we train a standard GCN (2 layers with 64 hidden units, 0.5 dropout rate, and ReLU) on the public benchmark Coauthor CS, with the 2-hop NC and a threshold of 0.4 to separate high-NC nodes from low-NC nodes.
@@
For nodes in a graph , the distinguishability is determined by the distribution . Under the transductive setting where all of the features and graph structure are accessible, different node subsets show different distinguishability from the distribution , where . If all nodes in the training set are trained by shared transformation weights, the test accuracy of different node subsets in the testing set reveals their distinguishability.
Observation 1.
Based on Proposition 2, we first calculate the NH values for nodes in the validation set and divide all nodes into ten groups according to the deciles of their NH values. Fig. 2 reports the average test accuracy of each group on 10 runs, along with the proportion of high-NC and low-NC nodes in each group. Consistent with previous studies, GCN achieves exceptionally high test accuracy within the group with the lowest NH, indicating that nodes in this group are not the least distinguishable. In conjunction with our NC metrics, this phenomenon can be easily explained by the negative correlation between test accuracy and the proportion of high-NC nodes. Specifically, the proportion of high-NC nodes in the group with NH of 0.0 is lower than that in the group with NH of 0.1, thereby resulting in a higher test accuracy. When the NH is 0.1 or higher, as it increases, the test accuracy rises as the proportion of high-NC nodes decreases. Such observation suggests that the NC metric can characterize the distinguishability of nodes to a certain extent. The results on more datasets are presented in Fig. 6.
@@
These phenomena support the argument that high-NC nodes confuse GCN. We describe this as nodes becoming confused when their surrounding neighbors hold diverse views. In such cases, nodes require clarification, that is, they are highly motivated to reconsider their raw features, discern messages from their noisy neighbors, and ultimately adopt targeted classification rules. Given the significant gap between high- and low-NC nodes, it is natural to optimize them separately to fit different learning processes. With our NCGCN model for separating the two groups of nodes during learning, high-NC nodes are distanced from the decision boundaries and are more evenly distributed within the clusters of their respective classes, as shown in Fig. 3. The test results show that the overall test accuracy improves to 95.67%, with 95.90% for the low-NC group and 82.47% for the high-NC group, yielding a significant boost in the latter.
@@
Node Confusion (NC) metric exhibits greater robustness compared to the Node Homophily (NH) metric. Unlike NH, which only focuses on consistency with the target node, NC takes into account the label distribution. Assuming a sufficiently large number of neighbors, NC ensures a continuous label perturbation, while in contrast, flipping the label of the target node can lead to an uncontrollable effect on NH. Moreover, compared to similarity-based metrics, it can be computed efficiently during model training. Compared to label-based metrics, it is free from the constraint of consistency with the target nodes, thereby allowing a larger range of neighbors to be counted.
@@ V Methodology
Based on the observations, we can conclude more fully that NC metric can measure the distinguishability of nodes, thus being more suitable for node separation. Due to the huge gap between these two groups of nodes, it is presumed that they necessitate distinct learning processes tailored to their respective levels of distinguishability. In this section, we propose the Neighborhood Confusion-guided Graph Convolutional Network (NCGCN). Fig. 4 shows the overall framework.
@@ V-A Neighborhood Confusion-guided Mask
In our framework, we separate the nodes into two groups, i.e., high- and low-NC nodes. Similar to [42, 45, 38], such absolute separation brings high computational efficiency. According to Eq. (3), the calculation of NC metrics relies on label information. However, it is not permissible to use all the ground-truth labels during training, so NC is calculated by pseudo labels. Specifically, we initialize NC values of all nodes as 0, and then update NC values each time the validation accuracy reaches a new high. A threshold is used to separate nodes into high- and low-NC groups, thus yielding the Neighborhood Confusion-guided Masks (NC mask) by
@@ V-C Model Training
V-D Plug-in Property and Efficiency
VI Evaluations
VI-A Experimental Settings
VI-A1 Datasets
We use 10 real-world benchmarks from [17, 46, 47], among which the first five (Pubmed, Photo, Computers, Coauthor CS and Coauthor Physics) are homophilous and the last five (Cora Full, Actor, Penn94, Chameleon and Squirrel) are heterophilous. We adopt de-duplicated version of the last two, namely Chameleon Filtered and Squirrel Filtered, as provided in [48]. For all datasets, we adopt dense data splitting as in [17], i.e., randomly split them into training, validation and testing sets with a proportion of 60%/20%/20%. We provide more details for these datasets in Appendix -A.
@@ VI-A1 Datasets
| NH Separ. | 90.88 0.56 | 94.54 0.42 | 89.71 0.53 | 95.75 0.41 | 98.06 0.40 | 66.17 0.73 | 39.03 1.63 | 78.52 0.51 | 39.24 6.64 | 31.81 5.08 |
@@ VI-A2 Competitors
We compare our methods with a total of 15 competitors in three categories, including classical methods such as two-layer MLP, GCN [9], SAGE [10] and GAT [11], spectral methods such as H2GCN [30], GPRGNN [17], FAGCN [21] and ACM-GCN [20], spatial methods such as CPGNN[43], WRGAT [37], GBKSage [26], HOG [22], GloGNN [49], CAGNN [40] and SNGNN [38]. The methods are the latest and tailor-made for the heterophily problem. We provide more details for these competitors in Appendix -B.

### v6

@@ Abstract
Graph neural networks (GNNs) have achieved remarkable advances in graph-oriented tasks. However, real-world graphs invariably contain a certain proportion of heterophilous nodes, challenging the homophily assumption of traditional GNNs and hindering their performance. Most existing studies continue to design generic models with shared weights between heterophilous and homophilous nodes. Despite the incorporation of high-order messages or multi-channel architectures, these efforts often fall short. A minority of studies attempt to train different node groups separately but suffer from inappropriate separation metrics and low efficiency. In this paper, we first propose a new metric, termed Neighborhood Confusion (NC), to facilitate a more reliable separation of nodes. We observe that node groups with different levels of NC values exhibit certain differences in intra-group accuracy and visualized embeddings. These pave the way for Neighborhood Confusion-guided Graph Convolutional Network (NCGCN), in which nodes are grouped by their NC values and accept intra-group weight sharing and message passing. Extensive experiments on both homophilous and heterophilous benchmarks demonstrate that our framework can effectively separate nodes and yield significant performance improvement compared to the latest methods. The source code is available in https://github.com/GISec-Team/NCGNN.
@@
Graph structured data can effectively model real-world interactive systems and has received considerable attention recently. Graph neural networks (GNNs) have been proven to be powerful in many graph-related applications, such as anomaly detection [1, 2, 3], fraud detection [4, 5, 6] and recommendation systems [7, 8, 9, 10]. Traditional GNNs [11, 12, 13] are adept at learning node representations in graphs where the vast majority of nodes are homophilous, i.e., nodes are primarily connected to nodes with same labels. However, real-world graphs often contain heterophilous nodes [14] and are known as heterophilous graphs when the proportion of these nodes is high, which challenges the homophily assumption of traditional GNNs and resulting in diminished performance [15, 16, 17]. In heterophilous graphs, the patterns of connections between nodes are usually complex and varied, rather than being limited to connections with similarly labeled nodes. This is referred to as the heterophily problem, and recent studies [18, 19] have highlighted that the key to addressing this problem is to focus on the diversity of nodes in the graph. Inspired by this, it’s natural to designing flexible model architectures for characterizing diverse nodes, whether homophilous or heterophilous.
@@
Most recent studies endeavor to design general models that still enable weight sharing between homophilous and heterophilous nodes. However, even with the incorporation of higher-order neighbors [20, 21, 22, 23] or multi-channel architectures [24, 25], these models fail to perform consistently well across all evaluation datasets. This may be due to two reasons: 1) these methods do not fundamentally alter the message-passing mechanism that essentially conforms to homophily assumption, thereby introducing noise [26]; 2) nodes within the same graph may exhibit significant behavioral differences due to distribution bias [27, 28]. For example, categorizing an interdisciplinary paper relies on more specific criteria compared to a paper confined to a single discipline [29]. Therefore, sharing weights universally across all nodes in a real-world graph is impractical, prompting the necessity for a metric to separate nodes with different characteristics. While numerous studies have explored the separation of neighbors [30, 31] or spectral signals [24, 25], the separation of target nodes or transformation weights, which we refer to as separated learning, remains under-explored. Only a minority of studies [14, 30] have focused on this separating scheme, but they suffer from incorrect separation metrics and low efficiency.
@@
Beginning with the classical Node Homophily (NH) metric [32] derived directly from the definition of homophily or heterophily, we analyze the bipartite graph scenario as depicted in Fig. 1. According to the node homophily metric defined in Eq. (1), the homophily of any node in the bipartite graph is zero. Nevertheless, graph convolutional network (GCN) can still perform well in such extremely heterophilous scenario, presenting a paradox [33]. This phenomenon has been widely observed in recent studies [33, 34, 24], yet they neither ease this anomaly directly with their models nor explore it on a wider range of real-world graphs. Moreover, analytical papers [35, 36, 37] reach the consensus that message passing with too many layers leads to over-smoothing or indistinguishability on node representations, highlighting the importance of distinguishability over consistency with central nodes for model performance. Therefore, we propose a new metric termed Neighborhood Confusion (NC) and detail its superiority over NH. Our proposed framework is guided by this metric for node separation, further validating the effectiveness of NC. Our contributions can be summarized as follows:
@@ II-A Heterophily Problem
Spatial methods allow for fine-grained adaptations of message passing and transformation rules [40]. Geom-GCN [32] and WRGAT [41] rewire graphs to aggregate neighbors with structural similarities in latent space. SNGNN [42] and GGCN [31] construct a similarity matrix, drawing upon embeddings from the preceding layer, to separate the dissimilar neighbors. These methods require computing new topologies or similarities and thus are time and memory consuming.
@@ IV Metrics and Observations
Heterophily is harmful to traditional GNNs, as different classes of nodes will be inappropriately mixed during message aggregation, leading to the indistinguishability of node representations [34]. However, GCN still performs well on bipartite graphs, which are completely heterophilous under the definition of NH, bringing paradox [33]. As shown in Fig. 1, means that NH metric considers the nodes in the bipartite graph to be extremely heterophilous. To alleviate this paradox, some new metrics have been proposed, such as aggregated similarity score [24], which is calculated by the similarity of intermediate representations, but with high computational complexity.
@@ IV-A Neighborhood Confusion
where is the set of label classes, normalizes NC to , and higher value indicates a greater degree of label confusion within target neighborhood. The receptive field of NC can be 1- or 2-hop by setting or , which is adaptively configured across different datasets.
@@
NC metric can outline the joint label distribution in the -hop ego-net of node , and can be approximated as a loose form of joint label distribution entropy in highly heterophilous scenarios, but with reduced computational complexity by focusing on the most frequent label in the ego-net.
@@
The joint probability that a target node and its neighbors in the -hop neighborhood have label can be denoted as , where is subject to the label distribution of neighbors. We will abbreviate the joint probability as for simplicity. Then the joint entropy of the label distribution around (and include) target node is:
Since NC is based on the most frequent label, we express it as a loose form of joint entropy. By defining and , we have:
Thus, as the label distribution in the ego-net becomes more uniform and diverse (i.e., the environment becomes more heterophilous), NC can be interpreted as an approximation of joint entropy . The range of is determined to be based on the Maximum Entropy Principle. Consequently, NC can be normalized within the interval by dividing . While joint entropy uses the full label distribution, NC only considers the most frequent label, reducing the complexity of the computation. ∎
@@ IV-B Observations
Rethinking the scenario of bipartite graph, as depicted in Fig. 1, suggests that our metric perceives the label distribution of node neighborhoods to have a low level of confusion, which is quite different from the result of NH. Then, we will further explain why NC surpasses NH in separating nodes from two observations. As an illustrative analysis, we train a standard GCN (2 layers with 64 hidden units, 0.5 dropout rate, and ReLU) on the public benchmark Coauthor CS [49], with the 2-hop NC and a threshold of 0.4 to separate high-NC nodes from low-NC nodes.
Observation 1.
We first calculate the NH values for nodes in the validation set and divide all nodes into ten groups according to the deciles of their NH values. Fig. 2 reports the average test accuracy of each group on 10 runs, along with the proportion of high-NC and low-NC nodes in each group. Consistent with previous studies, GCN achieves exceptionally high test accuracy within the group with the lowest NH, indicating that nodes in this group are not the least distinguishable. In conjunction with our NC metrics, this phenomenon can be easily explained by the negative correlation between test accuracy and the proportion of high-NC nodes. Specifically, the proportion of high-NC nodes in the group with NH of 0.0 is lower than that in the group with NH of 0.1, thereby resulting in a higher test accuracy. When the NH is 0.1 or higher, as it increases, the test accuracy rises as the proportion of high-NC nodes decreases. Such observation suggests that the NC metric can characterize the distinguishability of nodes to a certain extent. The results on more benchmark datasets are presented in Fig. 6.
@@
NC metric not only measures the distinguishability of nodes in a graph but also provides insight into the classification uncertainty for each node. Nodes with higher NC values are generally more challenging to classify due to the heterophily of their neighborhoods. The classification error rate for such nodes can be theoretically bounded using NC.
Proof.
To understand how neighborhood confusion impacts node classification, we first turn to conditional entropy , which represents the expected uncertainty in the target node’s label given the labels of its neighbors . In other works, reflects how much confusion remains about the target node’s label even after knowing its neighborhood’s labels. The conditional entropy can be formulated as follows:
where represents the conditional probability that node has label given its neighborhood labels .
Assuming that the label probability of the target node is determined by its neighborhood label distribution, we have:
For the most frequent label in the neighborhood, we have:
Let . The remaining labels () occupy the proportion , and for simplicity, we assume they are uniformly distributed, such that each of these labels has a conditional probability as follows:
Substituting the above conditional probabilities into the definition of conditional entropy, we obtain:
When the neighborhood size is sufficiently large, we can approximate . This allows us to establish the connection between conditional entropy and NC.
According to the definition of NC, we have:
Then the conditional probability of the most frequent label is:
Here we consider the base of the logarithm to be 2. Then, substiituting Eq. (12) into Eq. (10) yields:
Viewing conditional entropy as a function of NC, it is easy to prove that conditional entropy is positively correlated with NC in the interval (refer to Appendix -A for more details). In other words, as the neighborhood becomes more confused (with a higher NC value), the conditional entropy increases, indicating greater uncertainty in predicting the label of the target node based on neighborhood information. Therefore, the NC metric can measure the classification difficulty (or distinguishability) of nodes in a graph.
Next, we apply Fano’s Inequality [50] to establish a theoretical lower bound on the classification error rate based on the conditional entropy as follows:
We can see that the lower bound function on the right side remains positively correlated with NC, i.e., an increase in NC will lead to a larger lower bound on the classification error rate, indicating that nodes with higher NC values experience greater classification uncertainty and pose challenges for accurate classification. Furthermore, the classification error rate can be theoretically bounded using NC. The curve of conditional entropy and classification error bound versus NC under different numbers of label categories is shown in Appendix -A. ∎
@@
These phenomena support the argument that high-NC nodes confuse GCN. We describe this as nodes becoming confused when their surrounding neighbors hold diverse views. In such cases, nodes require clarification, that is, they are highly motivated to reconsider their raw features, discern messages from their noisy neighbors, and ultimately adopt targeted classification rules. Given the significant gap between high- and low-NC nodes, it is natural to optimize them separately to fit different learning processes. With our NCGCN model for separating the two groups of nodes during learning, high-NC nodes are distanced from the decision boundaries and are more evenly distributed within the clusters of their respective classes, as shown in Fig. 3(c). The test results show that the overall test accuracy improves to 95.67%, with 95.90% for the low-NC group and 82.47% for the high-NC group, yielding a significant boost in the latter.
@@
Node Confusion (NC) metric exhibits greater robustness compared to the Node Homophily (NH) metric. Unlike NH, which focuses more on the label consistency between the target node and its neighborhood, NC primarily considers the number of high-frequency labels. Assuming a sufficiently large number of neighbors, NC tends to be more stable when label perturbations occur in the neighborhood, whereas NH is more susceptible to uncontrollable effect, particularly when the label of the target node is disturbed. Moreover, compared to similarity-based metrics, NC can be computed efficiently during model training. Compared to label-based metrics, it is free from the constraint of consistency with the target nodes, thereby allowing a larger range of neighbors to be counted. Compared to difficulty metrics, NC does not require temporal tracking across multiple epochs, allowing for immediate node differentiation within a single training step. Compared to predictability metrics, NC encompasses all nodes, including those with high complexity, rather than excluding low-predictability or out-of-distribution samples, enabling a more comprehensive approach to node classification.
@@ V Methodology
Based on the observations, we can conclude that NC metric can measure the distinguishability of nodes, thus being more suitable for node separation. Due to the huge gap between these two groups of nodes, it is presumed that they necessitate distinct learning processes tailored to their respective levels of distinguishability. In this section, we propose the Neighborhood Confusion-guided Graph Convolutional Network (NCGCN). Fig. 4 shows the overall framework.
@@ V-A Neighborhood Confusion-guided Mask
In our framework, we separate the nodes into two groups, i.e., high- and low-NC nodes. Similar to [46, 51, 42], such absolute separation brings high computational efficiency. According to Eq. (3), the calculation of NC metrics relies on label information. However, it is not permissible to use all the ground-truth labels during training, so NC is calculated using pseudo labels, i.e., the model’s predicted labels for the nodes are used in place of the ground-truth labels for metric computation. Specifically, we initialize NC values of all nodes as 0, and then update NC values each time the validation accuracy reaches a new high. A threshold is used to separate nodes into high- and low-NC groups, thus yielding the Neighborhood Confusion-guided Masks (NC mask) by
@@ V-C Model Training
NH Separ. | 90.88 0.56 | 94.54 0.42 | 89.71 0.53 | 95.75 0.41 | 98.06 0.40 | 66.17 0.73 | 39.03 1.63 | 78.52 0.51 | 39.24 6.64 | 31.81 5.08 | ||||||||
[…]
V-D Plug-in Property and Efficiency
VI Evaluations
VI-A Experimental Settings
VI-A1 Datasets
We use 10 real-world benchmarks from [20, 49, 52], among which the first five (Pubmed, Photo, Computers, Coauthor CS and Coauthor Physics) are homophilous and the last five (Cora Full, Actor, Penn94, Chameleon and Squirrel) are heterophilous. Note that such division is based on the criterion of . We adopt de-duplicated version of the last two, namely Chameleon Filtered and Squirrel Filtered, as provided in [53]. For all datasets, we adopt dense data splitting as in [20], i.e., randomly split them into training, validation and testing sets with a proportion of 60%/20%/20%. We provide more details for these datasets in Appendix -B.
@@ VI-A2 Competitors
We compare our methods with a total of 15 competitors in three categories, including classical methods such as two-layer MLP, GCN [11], SAGE [12] and GAT [13], spectral methods such as H2GCN [34], GPRGNN [20], FAGCN [25] and ACM-GCN [24], spatial methods such as CPGNN[47], WRGAT [41], GBKSage [30], HOG [26], GloGNN [54], CAGNN [44] and SNGNN [42]. The methods are the latest and tailor-made for the heterophily problem. We provide more details for these competitors in Appendix -C.